Jim has $10000 to invest. He invests x dollars in tax-free municipal bonds yielding 8.5%. He invests 3 times the amount he invested in the bonds in certificates of deposit earning 9%. Jim puts the remainder of the money in a savings account at an interest rate of 5.25%.
A) express Jim's earned interest as a function of x.
B) How much total interest does Jim earn if he invests $3000 in certificates of deposits?
Please help with these questions, I am very confused.
Thank you very much.
Tony

Answer by stanbon(75887)

Jim has $10000 to invest. He invests x dollars in tax-free municipal bonds yielding 8.5%. He invests 3 times the amount he invested in the bonds in certificates of deposit earning 9%. Jim puts the remainder of the money in a savings account at an interest rate of 5.25%.
A) express Jim's earned interest as a function of x.
Int = 0.085x + 0.09(3x)
-------------------------------
B) How much total interest does Jim earn if he invests $3000 in certificates of deposits?
Int = 0.085*1000 + 0.09(3000) = 85 + 270 = $355.00
-------------------------------
Cheers,
Stan H.
==============